For Questions 1-15, consider a competitive market for a good where the demand curve is determined...

  1. For Questions 1-15, consider a competitive market for a good where the demand curve is determined by:

    the demand function: P = 5+-1*Qd and

    the supply curve is determined by the supply function: P = 0.5*Qs.

    Where P stands for Price, QD is quantity demanded and QS is quantity supplied.

    What is the quantity demanded of the good when the price level is P = $4?

QUESTION 2

  1. What is the quantity supplied of the good when the price level is P = $4?

QUESTION 3

  1. At P=$4 there is:

    a. Competitive equilibrium in the market;

    b. Excess supply in the market;

    c. Excess demand in the market;

    d. Rationing in the market;

QUESTION 4

  1. What is the value of the price elasticity of demand if the price of the good changes from Po=$4 to Pi=$5?

    NOTE: Express your answer in NEGATIVE terms.

QUESTION 5

  1. What is the value of the price elasticity of supply if the price of the good changes from Po=$4 to Pi=$5?

QUESTION 6

  1. What is the equilibrium quantity level for the good in the competitive market?

QUESTION 7

  1. What is the equilibrium price level for the good in the competitive market?

QUESTION 8

  1. What is the consumer surplus in the competitive market?

QUESTION 9

  1. What is the producer surplus in the competitive market?

QUESTION 10

  1. What is the total surplus in the competitive market?

QUESTION 11

  1. For Questions 11-15, additionally assume a market intervention of the form of per unit $2 tax on the consumption of the good.

    How many units of the good are sold in the market at equilibrium considering this market intervention?

QUESTION 12

  1. How much are consumers going to pay per unit of the good under this market intervention?

QUESTION 13

  1. How much is the per unit amount that producers will receive as payment under this market intervention?

  

QUESTION 14

  1. How much are the tax revenues under this market intervention?

QUESTION 15

  1. How much is the Dead Weight Loss under this market intervention?
Answer #1

AnsWER Demand - supply P-S-Qd. Qela 5-p. Pa 0.50s. Qs = 21 when p=$4 Quantity demand Qdo - 5-4 I unds. ( when p= 84 suppliedelasticity 1.00 1.AP 4004 -4 So elasticity of demand – 4. (5) Por or $4 & Piz 85 Dir 10 1. 405 - lordx100 Kloro OX & 25000 7.P= 2P 221167) Q- 3.34 P= 1.67 pnte. $1.67 at evilibrium above the equilibrium (8 consumer pare Surplus - and below It the isIt 10) bele is the area price and above the in the area below the equilibrium supply cerve is called producer Suoples 1x (167A at equilibrium zunits of goods are sold. (12) The consumer going to pay is $3. under this mortet 43 producer will intervent- IX (1.67-1) X (3.34-21 box o 64x +39 - v 4489 - 8.0.45 so Deadweight loss is - - 0.89 +0:45 $134